An inductor has a 53.6-? reactance when connected to a 60.0-Hz source. The inductor is removed and then connected to a 51.0-Hz source that produces a 80.0-V rms voltage. What is the maximum current in the inductor?
Answer #1

wL = 53.6

=> 60*2*3.14*L = 53.6

=> L = 0.14225053078556263269639 H

max current = 80*sqrt(2)/wL = (80*1.414213562373095)/(51*2*3.14*0.14225053078556263269639) = 2.483254718828964 A answer